can you please help me with {3a+b=3
{2a-5b=-15

Answer by kikerkins(17)

{3a+b=3
{2a-5b=-15
multiply the top 1 by 5
15a+ 5b=15
2a-5b=-15
17a=0
a=0
3(0)+b=3
b=3